# Continual Test-time Adaptation for End-to-end Speech Recognition on Noisy Speech

### v1

@@ 4.4.1 Single Domain
AA AC BA CM GS MU NB SD TP VC CHiME-3 Source model 40.6 27.7 66.9 49.7 75.6 51.4 120.1 19.4 25.8 49.7 30.0 Non-continual SUTA 30.6 17.4 53.7 38.7 54.5 39.0 112.3 15.0 17.4 39.3 23.3 SGEM 30.9 17.8 54.5 39.2 56.3 39.2 113.0 14.9 17.5 40.3 23.5 Continual CSUTA 39.8 22.6 63.4 53.4 58.4 54.7 68.1 23.2 23.0 50.9 27.6 AWMC 31.6 18.0 61.6 37.7 48.5 36.2 131.9 17.0 18.0 36.1 22.4 Fast-slow DSUTA 25.9 15.4 33.2 33.5 37.0 28.4 36.3 15.5 15.6 29.9 21.7
Table 1: WER(%) of different TTA methods on CHiME-3 and LibriSpeech-C with 10 types of noises. Reported WER is averaged over 3 runs.
We compare TTA performance on 11 domains by word error rate (WER) in Table 1. DSUTA shows significant improvement compared to the baseline methods. It outperforms both non-continual and continual baseline methods by a large margin, except for the SD domain, where it still achieves a 15.5% WER, close to SGEM’s performance (14.9%). Notably, on the NB domain, DSUTA achieves a 36.3% WER compared to SUTA, which has a WER greater than 100%, demonstrating the effectiveness of our method.
The key success factor of DSUTA is its ability to leverage learned knowledge from past samples. Figure 4 plots the WER difference compared to the pre-trained model on the CM domain over time. We compare three methods: SUTA with $N=10$, DSUTA with $(M,N)=(5,10)$, and DSUTA with $(M,N)=(5,0)$, i.e., the learned $\phi_{t}$ itself. The WER of $\phi_{t}$ is lower than that of the pre-trained model, and DSUTA with 10-step adaptation outperforms SUTA with 10-step adaptation. In other words, DSUTA adaptation has a "better start" compared to non-continual TTA methods due to the learned knowledge, resulting in superior performance. Furthermore, DSUTA is more efficient in adaptation steps than SUTA, requiring fewer steps to achieve better performance. Appendix A.2 provides more discussion on efficiency.
@@
Table 1 also compares other continual TTA methods. Naive continual training, such as CSUTA, results in unsatisfactory performance and is sometimes even worse than the original pre-trained model due to its instability. Although AWMC is designed to increase stability, its performance sometimes lags behind SUTA, particularly in cases where the original pre-trained model has an extremely high error rate(BA, NB). This is not surprising since AWMC relies on a pseudo-label approach. In contrast, DSUTA uses mini-batch gradient descent to enhance stability without the use of pseudo labels. Furthermore, the fast-slow approach allows DSUTA to inherit SUTA’s ability to better fit a single utterance, improving overall performance while avoiding the meta-parameters overfitting.
4.4.2 Time-varying Domains
@@
MD-Easy MD-Hard MD-Long Source model 32.7 74.6 61.0 Non-continual SUTA 24.0 60.4 53.3 SGEM 25.0 61.0 53.4 Continual CSUTA 37.3 83.6 100.3 AWMC 25.8 66.1 60.6 Fast-slow DSUTA 24.0 45.6 43.2 w/ Dynamic reset 22.7 39.8 35.8 w/ Fixed reset 22.8 49.4 45.2 w/ Oracle boundary 21.7 36.9 39.5
@@
Table 2 summarizes the results. DSUTA is comparable to or better than other baseline methods, and applying Dynamic reset further boosts the performance. Since we set DSUTA with fewer adaptation steps, our proposed method is both better and faster than SUTA in the multi-domain scenario [^7] See Appendix A.2 for detailed discussion on efficiency..
@@ 5.1 Why Choosing Averaged LII as an Indicator?
Figure 5: Distributions of averaged LII (over 5 samples) from the GS domain and non-GS domains.
Figure 6: Distributions of other possible $f_{D}$. (a): original LII, (b): averaged LII for 20 samples, (c): without subtraction of pre-trained model loss, and (d): with the adapted parameters.
Figure 6 visualizes the distributions of other possible choices of the indicator. Figure 6a, b shows the distribution of averaged LII over 1 sample (i.e., the original LII) and 20 samples, respectively. Using a single sample is not sufficient to distinguish the distributions, while considering more samples makes the detection more accurate. Figure 6c illustrates the case without subtracting the loss from the pre-trained model, namely $\mathcal{L}(\phi_{\mathcal{D}},x_{t})$. The distributions are not well separated. In Figure 6d, we also tried using the parameters after adaptation $A$ instead of the meta-parameters, namely
@@ 6 Conclusion
Different Types of End-to-End ASR Models: This work follows SUTA with a CTC-based ASR model, but there are different kinds of end-to-end ASR models available. As shown in Kim et al. (2023), entropy minimization-based TTA methods can be extended to other end-to-end ASR models. We encourage future research to extend our DSUTA method to these other end-to-end ASR models.
@@ A.1 Different noise levels
From Table 1 and Table 2, we observe a trend that DSUTA has a larger advantage over other methods under severe domain shift where the pre-trained model performs poorly. To investigate how different levels of domain shift affect the proposed method, we compare the pre-trained model, SUTA, and DSUTA with noise levels of 0dB, 5dB, and 10dB on the AC, SD, and TP domains from LibriSpeech-C, which are the top 3 well-performing domains for the pre-trained model. We set $N=5$ for both SUTA and DSUTA. Table 4 summarizes the results.
The results show that DSUTA is more effective under severe corruption. As the noise level decreases, although DSUTA outperforms the pre-trained model, SUTA becomes better than DSUTA. We hypothesize that while DSUTA is quite effective on noisy speech, its performance gain over the non-continual version (SUTA) is limited on relatively clean speech. Improving DSUTA’s performance over SUTA on clean speech remains an area for future work.
@@
A.2 Discussion on Efficiency
DSUTA is more efficient in adaptation steps than SUTA. Figure 7 compares SUTA and DSUTA on 10 domains of LibriSpeech-C under different adaptation steps $N=0,1,3,5,10$. DSUTA can use fewer adaptation steps to achieve better performance than SUTA with more adaptation steps.
Figure 7: WER(%) of different number of adaptation steps on 10 noise domains of LibriSpeech-C.
To assess the efficiency of different TTA methods, we run them on MD-Long and compare the required forward/backward steps in Table 5 and the runtime in Table 6. CSUTA is excluded due to its poor performance. We follow the hyperparameter settings described in Section 4.4.2. All experiments were conducted on an Nvidia GeForce RTX 3080Ti GPU. Note that the results are for reference only, as values can slightly differ depending on the implementation.
DSUTA is more efficient in the adaptation step and overall faster than SUTA, SGEM, and AWMC. Although adding the dynamic reset strategy slightly increases runtime, it remains faster overall. We conclude that our method is not only superior in performance but also more efficient than existing approaches.
#Forward #Backward
Non-continual
SUTA 100000 100000
SGEM 100000 100000
Continual
AWMC 300000 100000
Fast-slow
DSUTA 52000 52000
w/ Dynamic reset 72000 52000
Table 5: Forward/backward steps comparison for different TTA methods on MD-Long.
Total Avg
Non-continual
SUTA 5040 0.080
SGEM 11620 0.186
Continual
AWMC 11704 0.187
Fast-slow
DSUTA 3885 0.062
w/ Dynamic reset 4149 0.066
Table 6: Runtime(s) comparison for different TTA methods on MD-Long. Avg is the averaged runtime(s) for a 1-second utterance. The result is averaged over 3 runs.
A.3 Hyper-parameter Tuning
@@
A.4 Generalization to Different Source ASR Models
@@
A.5 SUTA’s Objective ($L_{suta}$)
@@
A.6 Pseudo Code for DSUTA with Dynamic Reset Strategy
Algorithm 2 describes the pseudo code of DSUTA with the dynamic reset strategy.

### v2

@@ 4.4.1 Single Domain
Method AA AC BA CM GS MU NB SD TP VC CHiME-3 Source model 40.6 27.7 66.9 49.7 75.6 51.4 120.1 19.4 25.8 49.7 30.0 Non-continual SUTA 30.6 17.4 53.7 38.7 54.5 39.0 112.3 15.0 17.4 39.3 23.3 SGEM 30.9 17.8 54.5 39.2 56.3 39.2 113.0 14.9 17.5 40.3 23.5 Continual CSUTA 39.8 22.6 63.4 53.4 58.4 54.7 68.1 23.2 23.0 50.9 27.6 AWMC 31.6 18.0 61.6 37.7 48.5 36.2 131.9 17.0 18.0 36.1 22.4 Fast-slow DSUTA 25.9 15.4 33.2 33.5 37.0 28.4 36.3 15.5 15.6 29.9 21.7
Table 1: WER (%) of different TTA methods on LS-C with 10 types of noises and CHiME-3. Reported WER is averaged over 3 runs.
We compare TTA performance on LS-C by Word Error Rate (WER) in Table 1. DSUTA shows significant improvement compared to the baseline methods. It outperforms both non-continual and continual baseline methods by a large margin, except for the SD domain, where it still achieves a 15.5% WER, close to SGEM’s performance (14.9%). Notably, on the NB domain, DSUTA achieves a 36.3% WER compared to SUTA, which has a WER greater than 100%, demonstrating the effectiveness of our method.
The key success factor of DSUTA is its ability to leverage learned knowledge from past samples. Figure 4 plots the WER difference compared to the pre-trained model on the CM domain over time. We compare three methods: SUTA with $N=10$, DSUTA with $(M,N)=(5,10)$, and DSUTA with $(M,N)=(5,0)$, i.e., the learned $\phi_{t}$ itself. The WER of $\phi_{t}$ is lower than that of the pre-trained model, and DSUTA with 10-step adaptation outperforms SUTA with 10-step adaptation. In other words, DSUTA adaptation has a “better start" compared to non-continual TTA methods due to the learned knowledge, resulting in superior performance.
@@
Table 1 also compares other continual TTA methods. Naive continual training, such as CSUTA, results in unsatisfactory performance and is sometimes even worse than the original pre-trained model due to its instability. Although AWMC is designed to increase stability, its performance sometimes lags behind SUTA, particularly in cases where the original pre-trained model has an extremely high error rate (BA and NB). This is not surprising since AWMC relies on a pseudo-label approach. In contrast, DSUTA uses mini-batch gradient descent to enhance stability without the use of pseudo labels. Furthermore, the fast-slow approach allows DSUTA to inherit SUTA’s ability to better fit a single utterance, improving overall performance while avoiding the meta-parameters overfitting.
4.4.2 Time-varying Multiple Domains
@@
Method MD-Easy MD-Hard MD-Long Source model 32.7 74.6 61.0 Non-continual SUTA 24.0 60.4 53.3 SGEM 25.0 61.0 53.4 Continual CSUTA 37.3 83.6 100.3 AWMC 25.8 66.1 60.6 Fast-slow DSUTA 24.0 45.6 43.2 w/ Dynamic reset 22.7 39.8 35.8 w/ Fixed reset 22.8 49.4 45.2 w/ Oracle boundary 21.7 36.9 39.5
@@
Table 2 summarizes the results. DSUTA is comparable to or better than other baseline methods, and applying Dynamic reset further boosts the performance. Since we set DSUTA with fewer adaptation steps, our proposed method is both better and faster than SUTA in the multi-domain scenario.
@@
Lastly, DSUTA demonstrates superior performance on real multi-domain noisy data, as shown in Table 1 column “CHiME-3". DSUTA achieves 21.7% WER, while the baseline SUTA and AWMC only yield 23.3% and 22.4% WER, respectively. This result further validates the proposed DSUTA can be generalized to real multi-domain noisy speech data stream.
@@ 5.1 Why Choosing Averaged LII as an Indicator?
Figure 5: Distributions of averaged LII (over 5 samples) from the GS domain (in) and non-GS domains (out).
Figure 6: Distributions of other possible indicators. (a): original LII, (b): averaged LII for 20 samples, (c): without subtraction of pre-trained model loss, and (d): with the adapted parameters.
Figure 6 visualizes the distributions of other possible choices of the indicator. Figure 6a, b shows the distribution of averaged LII over 1 sample (i.e., the original LII) and 20 samples, respectively. Using a single sample is not sufficient to distinguish the distributions while considering more samples makes the detection more accurate. Figure 6c illustrates the case without subtracting the loss from the pre-trained model, namely $\mathcal{L}(\phi_{\mathcal{D}},x_{t})$. The distributions are not well separated. In Figure 6d, we also tried using the parameters after adaptation $A$ instead of the meta-parameters, namely
@@ 5.2 Different Domain Transition Rates
Method Steps Runtime (s) #Forward #Backward Total Avg Non-continual SUTA 100000 100000 5040 0.080 SGEM 100000 100000 11620 0.186 Continual AWMC 300000 100000 11704 0.187 Fast-slow DSUTA 52000 52000 3885 0.062 w/ Dynamic reset 72000 52000 4149 0.066
Table 4: Comparison of Forward/Backward steps and Runtime for different TTA methods on MD-Long. Avg is the averaged runtime (s) for a 1-second utterance. The result is averaged over 3 runs.
5.3 Efficiency of the Proposed Method
DSUTA is more efficient in adaptation steps than SUTA. Appendix Figure 7 compares SUTA and DSUTA on 10 domains of LS-C under different adaptation steps $N=0,1,3,5,10$. DSUTA can use fewer adaptation steps to achieve better performance than SUTA with more adaptation steps.
To assess the efficiency of different TTA methods, we run them on MD-Long and compare the required forward/backward steps and runtime in Table 4. CSUTA is excluded due to its poor performance. We follow the hyperparameter settings described in Section 4.4.2. All experiments were conducted on an Nvidia GeForce RTX 3080Ti GPU. Note that the results are for reference only, as values can slightly differ depending on the implementation. DSUTA is more efficient in the adaptation step and overall faster than SUTA, SGEM, and AWMC. Although adding the dynamic reset strategy slightly increases runtime, it remains faster overall. In conclusion, our method is not only superior in performance but also more efficient than existing approaches.
5.4 Resets Frequency and Occurrence
We propose a dynamic model reset strategy to detect domain shifts, improving both performance and efficiency. However, the frequency of model resets and their positions within the data stream remain unclear. In Table 5, we present the reset timings of our method across three runs. The oracle boundaries occur at steps 500, 1000, 1500, and 2000. The results indicate that the reset timings are close to, but not exactly aligned with, the oracle boundaries.
3 runs Automatic reset step MD-Easy-1 540, 1530, 2010 MD-Easy-2 565, 1635, 2025 MD-Easy-3 560, 1530, 2010 MD-Hard-1 155, 555, 790, 1045 MD-Hard-2 510, 1510, 2010 MD-Hard-3 155, 510, 1165, 1510, 2010
Table 5: Reset times and Automatic reset steps for MD-Easy and MD-Hard tasks over 3 runs. The ground truth task boundaries at steps equal to 500, 1000, 1500, and 2000.
@@ 6 Conclusion
Different Types of End-to-End ASR Models: This work follows SUTA with a CTC-based ASR model, but there are different kinds of end-to-end ASR models available. As shown in (Kim et al., 2023), entropy minimization-based TTA methods can be extended to other end-to-end ASR models. We encourage future research to extend our DSUTA method to these other end-to-end ASR models.
@@
Acknowledgments
The authors want to thank the reviewers for their insightful comments and suggestions during the rebuttal period. Guan-Ting Lin is supported by the NTU GICE and the NTU Ph.D. scholarship.
@@ A.1 Different noise levels
From Table 1 and Table 2, we observe a trend that DSUTA has a larger advantage over other methods under severe domain shift where the pre-trained model performs poorly. To investigate how different levels of domain shift affect the proposed method, we compare the pre-trained model, SUTA, and DSUTA with noise levels of 0dB, 5dB, and 10dB on the AC, SD, and TP domains from LS-C, which are the top 3 well-performing domains for the pre-trained model. We set $N=5$ for both SUTA and DSUTA. Table 6 summarizes the results.
The results show that DSUTA is more effective under severe corruption. As the noise level decreases, although DSUTA outperforms the pre-trained model, SUTA becomes better than DSUTA. We hypothesize that while DSUTA is quite effective in noisy speech, its performance gain over the non-continual version (SUTA) is limited to relatively clean speech. Improving DSUTA’s performance over SUTA on clean speech remains an area for future work.
@@
Figure 7: WER (%) of different number of adaptation steps on 10 noise domains of LS-C.
A.2 Hyper-parameter Tuning
@@
A.3 Generalization to Different Source ASR Models
@@
A.4 Objective of SUTA ($L_{suta}$)